OpenStudy (anonymous):

The interval of convergence of the series [(3^n(x-2)^n)/ (n!)] is from n=0 to infinity.

OpenStudy (anonymous):

\[ \sum_{n=0}^{\infty} [(3^n(x-2)^n/ n!)]\]

OpenStudy (anonymous):

To be clear numerator: 3^n(x-2)^n Denominator: n!

OpenStudy (amistre64):

nice :)

OpenStudy (anonymous):

im so bad with series questions...i dont even know where to start with some of them

OpenStudy (amistre64):

i recall this vaguely, so ill have to resort to pauls workings of it

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/CalcII/PowerSeries.aspx if you wanna make sure i read it right ;)

OpenStudy (amistre64):

it looks like we take the ratio like in a ratio test

OpenStudy (anonymous):

yeah i rmbr my teaching saying if the series has a factorial and a power u use the ratio test

OpenStudy (amistre64):

\[\sum_{n=0}^{\infty} \frac{3^n(x-2)^n}{n!}\] \[\lim_{n\to inf} \frac{3^n(x-2)^n}{n!} \frac{(n-1)!}{3^{(n-1)}(x-2)^{(n-1)}}\] \[\lim_{n\to inf} \frac{3(x-2)}{n}\]

OpenStudy (amistre64):

as n get large; the value goes to zero

OpenStudy (anonymous):

i have a qustion

OpenStudy (amistre64):

which I believe means that it is convergent everywhere

OpenStudy (anonymous):

i thought for the ratio test ur supppsed to add 1 to n not subtract

OpenStudy (anonymous):

so like it shud be 3^(n+1)* (x-2)^n+1

OpenStudy (amistre64):

the ration test is of the form\[\frac{A_{now}}{A_{past}}\]or\[\frac{A_{future}}{A_{now}}\]

OpenStudy (amistre64):

the idea being that your ratio is a proportion of 2 terms with the top being the one right after the bottom

OpenStudy (amistre64):

so either n+1 the left part; or n-1 the right part

OpenStudy (anonymous):

wat do u mean left and right part

OpenStudy (anonymous):

srry my teacher never talked about ratio test like that

OpenStudy (amistre64):

\[\lim_{n\to inf} \frac{3^{(n+1)}(x-2)^{(n+1)}}{(n+1)!} \frac{n!}{3^{n}(x-2)^{n}}\] \[\lim_{n\to inf} \frac{3(x-2)}{n+1}\] same results over all

OpenStudy (amistre64):

since our ratio is a fraction over a fraction ..... \[\cfrac{\frac{a}{b}}{\frac{c}{d}}\to\cfrac{\frac{a}{b}}{\frac{c}{d}}*\cfrac{\frac{d}{c}}{\frac{d}{c}}= \frac{a}{b}\frac{d}{c}\]

OpenStudy (amistre64):

we end up with a left and a right part to this

OpenStudy (anonymous):

so now how do we find the interval of converegence since that is wat the question is asking

OpenStudy (amistre64):

well, we view the results of our limit ratio

OpenStudy (amistre64):

its zero; which tells us that the interval on which this converges is from -inf to inf

OpenStudy (amistre64):

example 4 of the link i posted gives a reasoning as to why i believe :)

OpenStudy (anonymous):

oh i see. thanks for helping. =)

OpenStudy (amistre64):

youre welcome